# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v2

@@ 3.1. SketchFDA: Sketch-based Estimation
An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix:
$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
The $\operatorname{sk}(\cdot)$ operator is linear and can be computed in $\mathcal{O}\left(l\cdot d\right)$ steps. Let $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$. The function
$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$.
In this approach, the lower-dimensional representation of the local drifts becomes $\sigma_{t}^{(k)}=\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}^{l\times m}$.
@@
Define:
$S_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
@@
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ with probability at least ($1-\delta$), i.e., we overestimate the model variance with probability at least ($1-\delta$), completing the proof.
@@ 3.2. LinearFDA: Linear Approximation
Although AMS sketches provide good estimates for variance, their dimension is in the several hundreds, and the communication cost of AllReduce on sketches, performed at each step, may be non-negligible. Therefore, we also examine a minimum-cost, ad-hoc estimation variant.
In this approach, $\mathbf{\sigma}_{t}^{(k)}$ is reduced to a scalar, that is, $\sigma_{t}^{(k)}=\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\in\mathbb{R}$ where $\xi\in\mathbb{R}^{d}$ is any unit vector.
@@
$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
@@
$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
@@
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof.
A random choice of $\xi$ is likely to perform poorly (force synchronization prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$. A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$. A heuristic choice is to take the difference of the models of the last two synchronizations, assuming that the current model drift is correlated with the model drift at the last synchronization. All workers can compute this heuristic choice of $\xi$ without communication cost:
$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}.$
@@ 3.3. Discussion
The main intuition for FDA is summarized in making the decision to synchronize dynamic, based on model variance during training. By Eq. 4, model variance is the difference between the average of squared (local) drifts $\frac{1}{K}\sum\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}$ and the squared average (global) drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$. When both are low, or when both are high, synchronizing the local models is not necessary.
When contrasting SketchFDA and LinearFDA, it should first be noted that both methods overestimate the true model variance. However, SketchFDA is probabilistically quite accurate. Our findings suggest that setting $l=5$ and $m=250$ (as defined in Section 3.1) yields excellent accuracy in practice. These values have been adopted in our experiments and are thus recommended. With these settings, the size of a sketch is $4k$ bytes, significantly smaller than the size of all our models.
Between the two variants of FDA, LinearFDA requires less computational effort and bandwidth to compute and to transmit its local state (the norm of a vector plus the dot product of two vectors, rather than a sketch). On the other hand, SketchFDA provides in practice a very accurate estimator of the true model variance which is expected to lead to fewer synchronizations.
An alternative point of view of FDA, relates to the well-known “bias vs. variance” trade-off (Goodfellow et al., 2016), which encapsulates the tension between under-fitting and over-fitting. Note that the term “variance” here refers to change to the learned model $\mathbf{w}$ when the training data changes, and is different from our model variance $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$.
@@ 4.1. Setup
Platform. We employ TensorFlow (Abadi et al., 2016), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We modified TensorFlow to include our FDA variants and all competitive algorithms. All relevant code, figures, and data of this study are available in https://github.com/miketheologitis/FedL-Sync-FDA.
@@
Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2021). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Nakandala and Kumar, 2022).
@@
Evaluation Methodology. Comparing DDL algorithms is not straightforward. For example, comparing DDL algorithms based on the average cost of a training epoch can be misleading, as it does not consider the effects on the trained model’s quality. To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in the Figures). Based on this definition, we focus on two performance metrics:
@@
    Communication cost, which is the total data (in bytes) transmitted by all workers. Translating this cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers. Its impact is larger in FL scenarios, where workers often use slower Wi-Fi connections.
@@
    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in the Figures) performed by each worker. Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the nodes. Its impact is larger for nodes with lower computational resources.
@@
Figure 2. LeNet-5 on MNIST. At Non-IID: Label ”0”, the samples of Label ”0” are assigned to few workers. At Non-IID: 60%, 60% of the dataset is sorted and allocated to workers, causing some workers to receive many samples from the same label
@@ 4.3. Results
Due to the extensive set of unique experiments (over $1000$), as detailed in Table 2, we leverage Kernel Density Estimation (KDE) plots (Waskom, 2021) to visualize the bivariate distribution of computation and communication costs incurred by each strategy for attaining the Accuracy Target. These KDE plots provide a high-level overview of the cost trade-off for training accurate models. As an illustrative example, Figure 2 depicts the strategies’ bivariate distribution for the LeNet-5 model trained on MNIST with different data heterogeneity setups. In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 2) that attained the Accuracy Target of $0.985$.
@@
Figure 6. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model.
@@
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 234, and 5 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 2-5, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 23,  4, and 5. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 3 for all three data heterogeneity settings. For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  4, and 5), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
FDA is resilient to data heterogeneity. In DDL, data heterogeneity is a prevalent challenge, reflecting the complexity of real-world applications where the IID assumption often does not hold. The ability of DDL algorithms to maintain consistent performance in the face of non-IID data is a critical metric for their effectiveness and adaptability. Our empirical investigation reveals the FDA methods’ noteworthy resilience in such heterogeneous environments. For LeNet-5 on MNIST, as illustrated in Figure 2, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label ”00”, $60\%$). Similarly, for VGG16* on MNIST, Figure 3 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label ”00”, Label ”$8$”); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
FDA generalizes better. The factors determining how well a DL algorithm performs are its ability to: (1) make the training accuracy high, and (2) make the gap between training and test accuracy small. These two factors correspond to the two central challenges in DL: underfitting and overfitting (Goodfellow et al., 2016). For DenseNet121 on CIFAR-10, with a test accuracy target of $0.8$, as illustrated in Figure 6, Synchronous and FedAvgM exhibit overfitting, with a noticeable discrepancy between training and test accuracy. In stark contrast, the FDA methods have an almost zero accuracy gap. Please note that LinearFDA and SketchFDA reach the test accuracy target of $0.8$ much earlier (at epochs 86 and 91, respectively). Turning our focus to DenseNet201 on CIFAR-10, with a test accuracy target of $0.78$, Synchronous again tends towards overfitting, while FedAvgM shows a slight improvement but still does not match the FDA methods, which continue to exhibit exceptional generalization capabilities, evidenced by a minimal training-test accuracy gap, as shown at Figure 6. Notably, given the necessity to fix hyper-parameters $\Theta$ and $K$ for the training accuracy plots, we selected two representative examples. The patterns of performance we highlighted are consistent across most of the conducted tests.

### v4

@@ 3.1. SketchFDA: Sketch-based Estimation
An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization and properties of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix:
$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\psi_{1}&\psi_{2}&\dots&\psi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
An estimate for squared-norm $\left\|\mathbf{v}\right\|_{2}^{2}$ is provided by the formula
$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\psi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
The quality of estimation depends on the size of the sketch. For chosen $\epsilon,\delta>0$, where sketch dimensions are given by $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$, we have the following probabilistic guarantee: with confidence at least $1-\delta$,
$\mathcal{M}_{2}(\operatorname{sk}(\mathbf{v}))\in(1\pm\epsilon)\left\|\mathbf{v}\right\|_{2}^{2}$
Notably, observe that the accuracy ($\epsilon$) and confidence ($1-\delta$) only depend on the size of the sketch and not on the dimensionality of vector $\mathbf{v}$.
Two crucial properties of the AMS sketch are that (a) it is a linear transformation, i.e., for $\alpha_{1},\alpha_{2}\in\mathbb{R}$ and $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{d}$,
$\operatorname{sk}(\alpha_{1}\mathbf{v}_{1}+\alpha_{2}\mathbf{v}_{2})=\alpha_{1}\operatorname{sk}(\mathbf{v}_{1})+\alpha_{2}\operatorname{sk}(\mathbf{v}_{2})$
and (b) can be computed efficiently in time $O(l\cdot d)$.
In the SketchFDA approach, the salient idea is to employ AMS sketches $\operatorname{sk}(\mathbf{u}_{t}^{(k)})\in\mathbb{R}^{l\times m}$ as a low-dimensional representation of the local drifts $\mathbf{u}_{t}^{(k)}$.
@@
Let $l=\mathcal{O}(\log\frac{1}{\delta})$ and $m=\mathcal{O}(\frac{1}{\epsilon^{2}})$. Define the local state as
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
and the approximation function as
@@
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ with probability at least ($1-\delta$), i.e., we overestimate the model variance with probability at least ($1-\delta$), completing the proof. ∎
In Section 3.3, we discuss the empirical basis for choosing the values of $l$ and $m$, and how they practically impact the quality of the sketch approximation.
@@ 3.2. LinearFDA: Linear Approximation
Although AMS sketches provide good estimates for variance, their dimension is in the several hundreds, and the communication cost of AllReduce on sketches, performed at each step, may be non-negligible. Therefore, we also introduce a low-cost, ad-hoc estimation variant.
In this approach, instead of an AMS sketch, each local state contains the scalar value $\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\rangle\in\mathbb{R}$, where $\xi\in\mathbb{R}^{d}$ is a unit vector, known to all workers.
@@
and the approximation function as
$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
@@
$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
@@
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof. ∎
An arbitrary choice of $\xi$ (e.g., a random vector) is likely to estimate $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ poorly; if $\mathbf{\xi}$ is uncorrelated to $\overline{\mathbf{u}}_{t}$, then $\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}$ will likely be close to zero. A heuristic choice that might be correlated to $\overline{\mathbf{u}}_{t}$ is the (normalized) value of $\overline{\mathbf{u}}_{t_{0}}$, the global drift vector right at the time of last synchronization. All nodes can compute it independently without extra communication, if they take the difference of the models of the last two synchronizations:
$\mathbf{\xi}=\frac{\overline{\mathbf{u}}_{t_{0}}}{\left\|\overline{\mathbf{u}}_{t_{0}}\right\|_{2}}=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}$
@@ 3.3. Discussion
FDA: Intuition. The main intuition for FDA is summarized in making the decision to synchronize dynamic, based on model variance during training. This metric is designed to capture the collective state of the training process. In what follows, we provide intuition on why this is the case. It is important to remember that the global model $\overline{\mathbf{w}}_{t}$ and, by extension, the global drift $\overline{\mathbf{u}}_{t}$, are ultimately what we care about and evaluate.
Model variance, as defined in Equation (4), is the difference between the average of the squared local drifts $\frac{1}{K}\sum\|\mathbf{u}_{t}^{(k)}\|_{2}^{2}$ and the squared global drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$. The first term reflects how far the individual worker models have moved–essentially, how much each worker has learned. The second term indicates how much of this learning is retained in the global model after aggregation.
The interplay between these two quantities is crucial. For example, when the local drifts are high but the global drift is low, the variance increases, signaling the need for synchronization. This scenario suggests that while individual workers have made significant progress (as indicated by high local drifts), this progress is not being effectively captured in the global model (indicated by the low global drift). In other words, the worker models have moved significantly, but the global model has remained relatively stationary in this high-dimensional space. This misalignment indicates that training is no longer progressing optimally, as the workers are moving towards disparate and conflicting local minima, making it crucial to synchronize and realign them. Conversely, when both the local and global drifts are either low or high, synchronization is not necessary, and the variance naturally remains low.
Neither the average of the local drifts nor the global drift alone provides a complete picture of the collective training progress. Relying solely on one or the other would lead to suboptimal synchronization decisions and likely prove ineffective. In FDA, it is the relationship between these quantities, as captured by the model variance, that offers valuable insights and guides the crucial decision of when to synchronize.
SketchFDA vs. LinearFDA: Both methods send the squared norm of the drift $\|\mathbf{u}_{t}^{(k)}\|_{2}^{2}$, but differ in the additional accompanying lower-dimensional representation they transmit (Figure 2):
  1. (1)
    SketchFDA: An AMS sketch of the local drift.
  2. (2)
    LinearFDA: The dot product of a vector and the local drift.
The key difference between these two variants lies in the fidelity of approximation of the model variance. While both methods conservatively overestimate the variance, SketchFDA provides a provably accurate estimation, which is expected to lead to fewer synchronizations. LinearFDA requires less computational effort and bandwidth to create and communicate the local states, but may overestimate variance by too much, causing unnecessary synchronizations.
Figure 2. SketchFDA & LinearFDA: Local State structure.
SketchFDA: Choice of $l$ and $m$. We empirically measured the approximation achieved with sketch dimensions of $l=5$ rows and $m=250$ columns (as defined in Section 3.1): these settings yield an error bound of $\epsilon\approx 6\%$ and a probabilistic confidence of $(1-\delta)\approx 95\%$. Based on our experiments, we have adopted these values in our experiments and recommend them. Using these values, the byte-size of a sketch is $l\cdot m\cdot 4\,\mathrm{bytes}=5\,\mathrm{kB}$, significantly smaller than the size of all our models. Sketches of smaller size could be used, albeit weakening the approximation of the variance. However, given that LinearFDA similarly weakens approximation and avoids using AMS sketches, in the interest of space we do not explore varying AMS sketch sizes in this paper.
FDA: Asynchronous Operation. As mentioned in Section 2, FDA can be readily modified to operate asynchronously. In this setup, one worker-node acts as a coordinator, aggregating local states and determining whether synchronization is needed each time a local state is received. This decision is based on the most recent local states from all workers. It is important to note that, since local states are small in size, asynchronous operation is unlikely to alleviate bandwidth issues. The primary advantage is that it allows training to continue even in the presence of stragglers. Asynchronous operation might also be beneficial in rare cases where the overhead of initializing communication dominates the actual transmission time.
@@ 4.1. Setup
Table 2 provides a comprehensive overview of our experiments. For each experiment, we detail the Neural Network (NN) architecture, its parameter count ($d$), and the dataset used for training. The table also specifies key hyper-parameters: the batch size ($b$), the number of workers ($K$), and the FDA-specific variance threshold ($\Theta$). Additionally, we indicate the chosen optimizer (as detailed in Section 3) and the training algorithms employed for each configuration.
Platform. We employ TensorFlow (Abadi et al., 2016), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We used TensorFlow to implement our FDA variants and all competitive algorithms. All relevant code, figures, and data of this study are available in https://github.com/miketheologitis/FedL-Sync-FDA.
Hardware & Infrastructure. We conducted our experiments on the ARIS High performance computing (HPC) environment [^2] https://www.hpc.grnet.gr/en/hardware-2/, utilizing a cluster of 44 GPU-accelerated worker-nodes. Each worker is equipped with two NVIDIA Tesla K40m GPUs and interconnected via an InfiniBand FDR14 network, providing up to 56 GB/s of bandwidth. Crucially, our evaluation remains agnostic to the underlying infrastructure of the specific workers.
@@
Lastly, we explore a transfer learning scenario on the dataset CIFAR-100 (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2021). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Nakandala and Kumar, 2022).
@@
Evaluation Methodology. Comparing DDL algorithms is not straightforward. For example, comparing DDL algorithms based on the average cost of a training epoch can be misleading, as it does not consider the effects on the trained model’s quality. To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in figures). Based on this definition, we focus on two performance metrics:
@@
    Communication cost, which is the total data (in bytes) transmitted by all workers. Notably, communication cost is unaffected by the training data volume since only model updates (when synchronizing) and local states (at each step), but not training data, are transmitted. Thus, the communication cost mainly depends on the complexity (number of parameters) of the used model. Translating the communication cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers and on the overhead of establishing and initializing communication. Its impact is larger in FL scenarios, where workers often use slower Wi-Fi connections.
@@
    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in figures) performed by each worker. Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the worker-nodes. Its impact is larger for workers with lower computational resources.
@@
Figure 3. LeNet-5 on MNIST. At Non-IID: Label "0", the samples of Label "0" are assigned to few workers. At Non-IID: 60%, 60% of the dataset is sorted and allocated to workers, causing some workers to receive many samples from the same label
@@ 4.3. Results
Due to the extensive set of unique experiments (over $1000$), as detailed in Table 2, we leverage Kernel Density Estimation (KDE) plots (Waskom, 2021) to visualize the bivariate distribution of computation and communication costs incurred by each strategy for attaining the Accuracy Target. These KDE plots provide a high-level overview of the cost trade-off for training accurate models. The varying levels of opacity in the filled areas of the KDE plots represent the density of the underlying data points: higher opacity indicates areas with a greater concentration of data, whereas lower opacity signifies less dense areas.
As an illustrative example, Figure 3 depicts the strategies’ bivariate distribution for the LeNet-5 model trained on MNIST with different data heterogeneity setups. In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 2) that attained the Accuracy Target of $0.985$. The observed high variance in the method’s distribution stems from the varying $K$ and $\Theta$ values. In subsequent subsections, we elucidate how these hyper-parameters influence the communication and computation costs.
@@
Figure 7. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (top), and $0.78$ (bottom). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 346, and 6 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 3-6, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 34,  6, and 6. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 4 for all three data heterogeneity settings. For a 0.001 increase in accuracy (effectively 10 misclassified testing images), FedAdam needs approximately 2-7$\times$ more communication and 3-7$\times$ more computation, respectively. This can be seen by comparing the figures at the left column of Figure 4 with the corresponding ones in the right column. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  6, and 6), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
FDA is resilient to data heterogeneity. In DDL, data heterogeneity is a prevalent challenge, reflecting the complexity of real-world applications where the IID assumption often does not hold. The ability of DDL algorithms to maintain consistent performance in the face of non-IID data is a critical metric for their effectiveness and adaptability. Our empirical investigation reveals the FDA methods’ noteworthy resilience in such heterogeneous environments. For LeNet-5 on MNIST, as illustrated in Figure 3, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label "00", $60\%$). Similarly, for VGG16* on MNIST, Figure 4 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label "00", Label "$8$"); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
FDA has a lower generalization gap. The factors determining how well a DL algorithm performs are its ability to: (1) make the training accuracy high, and (2) make the gap between training and test accuracy small. These two factors correspond to the two central challenges in DL: underfitting and overfitting (Goodfellow et al., 2016). For DenseNet121 on CIFAR-10, with a test accuracy target of $0.8$, as illustrated in Figure 7, Synchronous and FedAvgM exhibit overfitting, with a noticeable discrepancy between training and test accuracy. In stark contrast, the FDA methods have an almost zero accuracy gap. Please note that LinearFDA and SketchFDA reach the test accuracy target of $0.8$ much earlier (at epochs 86 and 91, respectively). Turning our focus to DenseNet201 on CIFAR-10, with a test accuracy target of $0.78$, Synchronous again tends towards overfitting, while FedAvgM shows a slight improvement but still does not match the FDA methods, which continue to exhibit exceptional generalization capabilities, evidenced by a minimal training-test accuracy gap, as shown at Figure 7. Notably, given the necessity to fix hyper-parameters $\Theta$ and $K$ for the training accuracy plots, we selected two representative examples. The patterns of performance we highlighted are consistent across most of the conducted tests.